# In this paper, we investigate the mixed norm estimates for the operator $T$associated with a dilated plane curve $(ut,u\gamma(t))$, defined by

### v2

@@ 3.1 Sufficiency of the region $(\frac{1}{p},\frac{1}{q})$ for all $r\in[1,\infty]$
We begin the proof with the following lemma.
@@
[…]
@@
for all multi-indices $\beta$ with $|\beta|=2$. Noting that $\partial_{\xi}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))=x-u(t_{0},\Gamma_{j}(t_{0}))$ and $|(t_{0},\Gamma_{j}(t_{0}))|\approx 1$, integration by parts (using Stein [46, Chapter VIII, Section 2, Proposition 5] and (2.4)) yields
@@
From [34, (2.44)], for any $\epsilon\in(0,\infty)$, we have
@@
[…]
  1. $\bullet$
    Decay bound at $(\frac{1}{2},\frac{1}{2})$: For $r\in[1,2]$, by Hölder’s inequality and (3.6),
    $\left\|F_{j,k}f\right\|_{L_{x}^{2}L_{u}^{r}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\left\|F_{j,k}f\right\|_{L_{x}^{2}L_{u}^{2}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{-\frac{1}{2}k}\|f\|_{L_{x}^{2}(\mathbb{R}^{2})}.$
    For $r\in[2,\infty]$, interpolation between (3.6) and (3.7) yields
    $\left\|F_{j,k}f\right\|_{L_{x}^{2}L_{u}^{r}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{-\frac{1}{r}k}\|f\|_{L_{x}^{2}(\mathbb{R}^{2})}.$
@@ 6. $\bullet$
Therefore, this completes the proof of the sufficiency of the region $(\frac{1}{p},\frac{1}{q})$ in Theorem 1.1 for all $r\in[1,\infty]$.
In this subsection, we establish the necessary conditions for Theorem 1.1. The condition $0\leq\frac{1}{q}$ is trivial, and the condition $\frac{1}{q}\leq\frac{1}{p}$ is a standard necessity following from a theorem of Hörmander [19]. We now demonstrate that, if
holds, then $(\frac{1}{p},\frac{1}{q})$ satisfies
$1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)\geq 0$
The proof is similar to [31]. Specifically, it suffices to show
where the implicit constant is independent of $i$, and $\{t_{i}\}_{i=1}^{\infty}$ strictly decreasing to 0.
Indeed, let $S_{t_{i}}:=[-t_{i},t_{i}]\times[-|\gamma(t_{i})|,|\gamma(t_{i})|]$. As shown in [31] that
$\delta t_{i}\chi_{\frac{S_{t_{i}}}{2}}(x)\leq T\chi_{S_{t_{i}}}(x,u)$
To show the sharpness of the other region of $(\frac{1}{p},\frac{1}{q})$, we first establish space-time estimates for Fourier integral operators, defined on the space of Schwartz functions $\mathcal{S}(\mathbb{R}^{n})$, of the form
@@ 3.2 The necessity of the region of $(\frac{1}{p},\frac{1}{q})$ for all $r\in[1,\infty]$
[…]
@@ 2. $\blacksquare$
Let $1\leq p\leq q\leq\infty$ and $s\in\mathbb{R}$. Let $P_{k}$ be the Littlewood-Paley operator defined in Section 111, and $\mathcal{F}$ be the Fourier integral operator defined in (3.10) with a symbol of order $\sigma$ satisfying the cinematic curvature condition, together with the special assumption that $\textrm{rank}~\partial^{2}_{\xi\xi}\phi(z,\xi)=n-1$ for $\xi\neq 0$. Then, there exists a positive constant $C_{(u)}$ such that
@@
The case $q=p\in(1,\infty)$ has been proved in Seeger, Sogge and Stein [41, Theorem 2.1]. Then, by interpolation, it suffices to prove
@@
Now, we suppose that there exists such $v$. Without loss of generality we may assume that $v=e_{1}$, otherwise we can choose new axes so that $v=e_{1}$. Since $\phi(z,\xi)$ is homogeneous of degree 111 in $\xi$,
@@
It is easy to see that, for any $i,j=1,2,\cdots,n$, $\partial^{2}_{\xi_{i}\xi_{j}}\phi(z,\xi)$ is homogeneous of degree 111 in $\xi$ and $\partial_{\xi_{i}}\phi(z,\xi)$ is homogeneous of degree 0 in $\xi$. Let $\phi_{\xi_{1}}(z,\xi^{\prime}):=\phi(z,\xi)$. Then, for all $|(\xi_{1},0)|\in\textrm{supp}~\bar{\psi}$, by (3.14) and (3.15), we can see that
@@
where $y:=(y_{1},y^{\prime})$. Therefore, from Stein [46, Chapter VIII, Proposition 6], there exists a constant $C$ such that the following identity is valid:
@@
We now turn to the proof of (3.12), which is based on some results obtained in Stein [46, Chapter IX, Section 4]. Let $\mathcal{F}$ be the Fourier integral operator with a symbol of order $-\frac{n-1}{2}$, it is enough to prove
@@
Next, a further dyadic division is needed. Roughly speaking, we will split each dyadic shell $\{\xi\in\mathbb{R}^{n}:2^{k-3}\leq|\xi|\leq 2^{k+3}\}$ into thin truncated cones of aperture roughly $2^{-\frac{k}{2}}$, and each such truncated cone is essentially an elongated rectangle whose major axis has length roughly $2^{k}$, while all the other sides have length roughly $2^{\frac{k}{2}}$. Then, we need roughly $2^{\frac{n-1}{2}k}$ such truncated cones to cover the shell $\{\xi\in\mathbb{R}^{n}:2^{k-3}\leq|\xi|\leq 2^{k+3}\}$ (see Stein [46, Chapter IX, Figure 1]). More precisely, for each $k\in\mathbb{N}$, we fix a collection $\{\xi_{k}^{\nu}\}_{\nu}\subset\mathbb{S}^{n-1}$ satisfy:
@@
As in Stein [46, Chapter IX, Section 4], for any $N\in\mathbb{N}_{0}$, we have
@@
where $(\cdot)_{1}$ denotes the component in the direction $\xi_{k}^{\nu}$ and $(\cdot)^{\prime}$ denotes the orthogonal component. Moreover, by the fact that $\mathcal{F}$ has a symbol of order $-\frac{n-1}{2}$, it follows that
@@
Let $n=2$, $1\leq p\leq q\leq\infty$, and $s\in\mathbb{R}$. Let $P_{k}$ be the Littlewood-Paley operator defined in Section 111, and $\mathcal{F}$ be a Fourier integral operator as in (3.10) with a symbol of order $\sigma$ satisfying the cinematic curvature condition, and assume $\textrm{rank}~\partial^{2}_{\xi\xi}\phi(z,\xi)=1$ for $\xi\neq 0$. Then, for $P_{k}$ defined as in Section 1 with $k\in\mathbb{N}$, there exists a constant $C>0$ such that
@@
[…]
@@
[…]
@@
where $N\in\mathbb{N}$ is large enough. Furthermore, from Mockenhaupt, Seeger and Sogge [35, P. 70], by the above non-degeneracy condition, we can choose a locally coordinates $z=(x,u)\in\mathbb{R}^{2}\times\mathbb{R}$ such that
@@
[…]
@@
From [31], we can see that the result in Lemma 3.3 is nearly sharp. Moreover, if we restrict our attention to the Littlewood-Paley operator $P_{k}$, Lemma 3.3 extends the $L_{x}^{2}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate obtained in Mockenhaupt, Seeger and Sogge [35, Theorem 3.1] or the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate with $q\geq 3p^{\prime}$ and $q\geq\frac{5}{3}p$ proved in Lee [26, Corollary 1.5] to an almost sharp $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate. Specifically, we have
@@
which follows from Seeger, Sogge and Stein [41, Theorem 3.1] by the fact that $F_{j,k}$ is a Fourier integral operator of order $-\frac{1}{2}$ satisfying the cinematic curvature condition. We remark that the non-degeneracy assumptions (2.2)–(2.3) and the cone condition (2.6) in Mockenhaupt, Seeger and Sogge [35, Theorem 3.1] are equivalent to the cinematic curvature condition.
On the other hand, we may see from [31] that the estimate (3.21) is an endpoint estimate, which implies the sharpness of the $L^{2}_{\textrm{comp}}(Y)\rightarrow L^{q}_{\textrm{loc}}(Z)$ local smoothing theorem [35, Theorem 3.1], and further implies the sharpness of the square function estimate [35, Theorem 3.2] by using Sobolev’s embedding theorem in the time variable $u$, at least in two-dimensional situation. Here, we also note that the results in [35, Theorems 3.1 and 3.2] actually are sharp for any $n$-dimensional setting by using the counterexample in Sogge [42, Section 6.2, P. 185].
We also require the following lemma.
Lemma 3.6.
Proof.
For $r=\infty$, the result is trivial. For $r\in[1,\infty)$, note that
Based on Lemmas 3.2, 3.3 and 3.6, we have the following proposition.

### v3

@@ 3.1 Sufficiency of the region $(\frac{1}{p},\frac{1}{q})$ for all $r\in[1,\infty]$
We begin the proof with the following two lemmas.
@@
$\left\{\begin{aligned} &\partial_{\xi_{1}\xi_{1}}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))=u\frac{1}{\xi_{2}}\frac{1}{\Gamma^{\prime\prime}_{j}(t_{0})};\\ &\partial_{\xi_{1}\xi_{2}}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))=-u\frac{\xi_{1}}{\xi_{2}^{2}}\frac{1}{\Gamma^{\prime\prime}_{j}(t_{0})};\\ &\partial_{\xi_{2}\xi_{1}}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))=-u\frac{\xi_{1}}{\xi_{2}^{2}}\frac{1}{\Gamma^{\prime\prime}_{j}(t_{0})};\\ &\partial_{\xi_{2}\xi_{2}}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))=u\frac{\xi_{1}^{2}}{\xi_{2}^{3}}\frac{1}{\Gamma^{\prime\prime}_{j}(t_{0})}.\end{aligned}\right$.
@@
for all multi-indices $\beta$ with $|\beta|=2$. Noting that $\partial_{\xi}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))=x-u(t_{0},\Gamma_{j}(t_{0}))$ and $|(t_{0},\Gamma_{j}(t_{0}))|\approx 1$, integration by parts (using Stein [47, Chapter VIII, Section 2, Proposition 5] and (2.4)) yields
@@
Lemma 3.2.
Proof.
For $r=\infty$, the result is trivial. For $r\in[1,\infty)$, note that
@@
From [35, (2.44)], for any $\epsilon\in(0,\infty)$, we have
@@
From Figure 1, for all $r\in[1,\infty]$, it suffices to obtain a constant bound $C$ or a small bound $2^{\epsilon_{1}k}$ at points $\{\text{O, A, B, C, D}\}$ for any $\epsilon_{1}\in(0,\infty)$, and a decay bound $2^{-\epsilon_{2}k}$ for $F_{j,k}$ at $(\frac{1}{4},\frac{1}{4})$ with some $\epsilon_{2}\in(0,\infty)$. More precisely, we first claim the following:
@@ 5. $\bullet$
Therefore, we reduce our proof to obtaining a decay bound at $(\frac{1}{4},\frac{1}{4})$. For any $r\in[1,\infty]$, by Hölder’s inequality, it is easy to see that
$\left\|F_{j,k}f\right\|_{L_{x}^{4}L_{u}^{r}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\left\|F_{j,k}f\right\|_{L_{x}^{4}L_{u}^{\infty}(\mathbb{R}^{2}\times\mathbb{R})}.$
Then, we just need to prove
(3.8) $\left\|F_{j,k}f\right\|_{L_{x}^{4}L_{u}^{\infty}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{\left(-\frac{1}{4}+\epsilon\right)k}\|f\|_{L_{x}^{4}(\mathbb{R}^{2})}$
for any $\epsilon\in(0,\infty)$. It is easy to see that the desired decay bound at $(\frac{1}{4},\frac{1}{4})$ can be obtained by letting $\epsilon\in(0,\frac{1}{4})$ in (3.8).
To obtain (3.8), we first establish the space-time estimates for Fourier integral operators, defined on the space of Schwartz functions $\mathcal{S}(\mathbb{R}^{n})$, of the form
@@
where $a(z,\xi)$ is a symbol of order $\sigma\in\mathbb{R}$. Here and hereafter, $z$ denotes a vector in $\mathbb{R}^{n}\times\mathbb{R}$ comprising the space-time variables $(x,u)$. Assume that $\textrm{supp}~a(\cdot,\xi)$ is contained in a fixed compact set and $\phi(z,\cdot)$ is homogeneous of degree 1. We say that $\mathcal{F}$ satisfies the cinematic curvature condition if the following hold:
@@ 2. $\blacksquare$
We now establish the following two lemmas, which will be used repeatedly in the remaining parts of this paper.
@@
Let $1\leq p\leq q\leq\infty$ and $s\in\mathbb{R}$. Let $P_{k}$ be the Littlewood-Paley operator defined in Section $1$, and $\mathcal{F}$ be the Fourier integral operator defined in (3.9) with a symbol of order $\sigma$ satisfying the cinematic curvature condition, together with the special assumption that $\textrm{rank}~\partial^{2}_{\xi\xi}\phi(z,\xi)=n-1$ for $\xi\neq 0$. Then, there exists a positive constant $C_{(u)}$ such that
@@
The case $q=p\in(1,\infty)$ has been proved in Seeger, Sogge and Stein [42, Theorem 2.1]. Then, by interpolation, it suffices to prove
@@
Now, we suppose that there exists such $v$. Without loss of generality we may assume that $v=e_{1}$, otherwise we can choose new axes so that $v=e_{1}$. Since $\phi(z,\xi)$ is homogeneous of degree $1$ in $\xi$,
@@
It is easy to see that, for any $i,j=1,2,\cdots,n$, $\partial^{2}_{\xi_{i}\xi_{j}}\phi(z,\xi)$ is homogeneous of degree $1$ in $\xi$ and $\partial_{\xi_{i}}\phi(z,\xi)$ is homogeneous of degree 0 in $\xi$. Let $\phi_{\xi_{1}}(z,\xi^{\prime}):=\phi(z,\xi)$. Then, for all $|(\xi_{1},0)|\in\textrm{supp}~\bar{\psi}$, by (3.13) and (3.14), we can see that
@@
where $y:=(y_{1},y^{\prime})$. Therefore, from Stein [47, Chapter VIII, Proposition 6], there exists a constant $C$ such that the following identity is valid:
@@
We now turn to the proof of (3.11), which is based on some results obtained in Stein [47, Chapter IX, Section 4]. Let $\mathcal{F}$ be the Fourier integral operator with a symbol of order $\frac{n-1}{2}$, it is enough to prove
@@
Next, a further dyadic division is needed. Roughly speaking, we will split each dyadic shell $\{\xi\in\mathbb{R}^{n}:2^{k-3}\leq|\xi|\leq 2^{k+3}\}$ into thin truncated cones of aperture roughly $2^{-\frac{k}{2}}$, and each such truncated cone is essentially an elongated rectangle whose major axis has length roughly $2^{k}$, while all the other sides have length roughly $2^{\frac{k}{2}}$. Then, we need roughly $2^{\frac{n-1}{2}k}$ such truncated cones to cover the shell $\{\xi\in\mathbb{R}^{n}:2^{k-3}\leq|\xi|\leq 2^{k+3}\}$ (see Stein [47, Chapter IX, Figure 1]). More precisely, for each $k\in\mathbb{N}$, we fix a collection $\{\xi_{k}^{\nu}\}_{\nu}\subset\mathbb{S}^{n-1}$ satisfy:
@@
As in Stein [47, Chapter IX, Section 4], for any $N\in\mathbb{N}_{0}$, we have
@@
where $(\cdot)_{1}$ denotes the component in the direction $\xi_{k}^{\nu}$ and $(\cdot)^{\prime}$ denotes the orthogonal component. Moreover, by the fact that $\mathcal{F}$ has a symbol of order $\frac{n-1}{2}$, it follows that
@@
Let $n=2$, $1\leq p\leq q\leq\infty$, and $s\in\mathbb{R}$. Let $P_{k}$ be the Littlewood-Paley operator defined in Section $1$, and $\mathcal{F}$ be a Fourier integral operator as in (3.9) with a symbol of order $\sigma$ satisfying the cinematic curvature condition, and assume $\textrm{rank}~\partial^{2}_{\xi\xi}\phi(z,\xi)=1$ for $\xi\neq 0$. Then, for $P_{k}$ defined as in Section 1 with $k\in\mathbb{N}$, there exists a constant $C>0$ such that
@@
The proofs of estimates $(\textrm{a})-(\textrm{d})$ follow directly from Lemma 3.3, since $C_{(u)}\lesssim 1$ for any $u\in[1,2]$. On the other hand, we want to show that Lee [26, Corollary 1.5] also implies the estimates $(\textrm{a})-(\textrm{b})$, where the following additional condition is needed:
@@
The estimate $(\textrm{c})$ can also be proved in the similar way as (3.10). Indeed, by the fact that $\phi(z,\xi)$ is homogeneous of degree $1$ in $\xi$, we can write (3.15) as
@@
where $N\in\mathbb{N}$ is large enough. Furthermore, from Mockenhaupt, Seeger and Sogge [36, P. 70], by the above non-degeneracy condition, we can choose a locally coordinates $z=(x,u)\in\mathbb{R}^{2}\times\mathbb{R}$ such that
@@
This, combined with the fact that $\textrm{supp}~a(\cdot,\xi)$ is contained in a fixed compact set, leads to the estimate $(\textrm{c})$.
The estimate $(\textrm{d})$ is a basic result about the Fourier integral operators $\mathcal{F}$, which can be proved by considering the pseudo-differential operator $\mathcal{F}^{\ast}\mathcal{F}$, we refer the reader to $\textrm{H}\ddot{\textrm{o}}\textrm{rmander}$ [20], Stein [47, Chapter IX, Section 3] or Sogge [43, Section 6.2] for more detailed discussions. The key estimate $(\textrm{e})$ follows from Gao, Liu, Miao and Xi [12, Theorem 1.4], which is a generalization of the local smoothing conjecture in $2+1$ dimensions obtained in Guth, Wang and Zhang [16, Theorem 1.2]. This finishes the proof. ∎
@@
From [31], we can see that the result in Lemma 3.4 is nearly sharp. Moreover, if we restrict our attention to the Littlewood-Paley operator $P_{k}$, Lemma 3.4 extends the $L_{x}^{2}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate obtained in Mockenhaupt, Seeger and Sogge [36, Theorem 3.1] or the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate with $q\geq 3p^{\prime}$ and $q\geq\frac{5}{3}p$ proved in Lee [26, Corollary 1.5] to an almost sharp $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate. Specifically, we have
@@
which follows from Seeger, Sogge and Stein [42, Theorem 3.1] by the fact that $F_{j,k}$ is a Fourier integral operator of order $\frac{1}{2}$ satisfying the cinematic curvature condition. We remark that the non-degeneracy assumptions (2.2)–(2.3) and the cone condition (2.6) in Mockenhaupt, Seeger and Sogge [36, Theorem 3.1] are equivalent to the cinematic curvature condition.
On the other hand, we may see from [31] that the estimate (3.20) is an endpoint estimate, which implies the sharpness of the $L^{2}_{\textrm{comp}}(Y)\rightarrow L^{q}_{\textrm{loc}}(Z)$ local smoothing theorem [36, Theorem 3.1], and further implies the sharpness of the square function estimate [36, Theorem 3.2] by using Sobolev’s embedding theorem in the time variable $u$, at least in two-dimensional situation. Here, we also note that the results in [36, Theorems 3.1 and 3.2] actually are sharp for any $n$-dimensional setting by using the counterexample in Sogge [43, Section 6.2, P. 185].
Now, we turn to the proof of (3.8). Notice that $F_{j,k}$ is a Fourier integral operator of order $\frac{1}{2}$ and satisfies (2.10), then the estimate $(\textrm{e})$ in Lemma 3.4 implies
(3.21) $\left\|F_{j,k}f\right\|_{L_{x}^{4}L^{4}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{(-\frac{1}{2}+\epsilon)k}\|f\|_{L_{x}^{4}(\mathbb{R}^{2})}$
for any $\epsilon\in(0,\infty)$. Similarly, we can obtain that
(3.22) $\left\|\partial_{u}(F_{j,k}f)\right\|_{L_{x}^{4}L^{4}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{(\frac{1}{2}+\epsilon)k}\|f\|_{L_{x}^{4}(\mathbb{R}^{2})}.$
Therefore, as an immediate consequence of (3.21) and (3.22), by Lemma 3.2, we deduce that
$\left\|F_{j,k}f\right\|_{L_{x}^{4}L^{\infty}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim\left\|F_{j,k}f\right\|^{\frac{3}{4}}_{L_{x}^{4}L^{4}_{u}(\mathbb{R}^{2}\times[1,2])}\left\|\partial_{u}(F_{j,k}f)\right\|^{\frac{1}{4}}_{L_{x}^{4}L^{4}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{(-\frac{1}{4}+\epsilon)k}\|f\|_{L_{x}^{4}(\mathbb{R}^{2})}$
for any $\epsilon\in(0,\infty)$, which is (3.8) as desired.
Putting things together, we complete the proof of the sufficiency of the region $(\frac{1}{p},\frac{1}{q})$ in Theorem 1.1 for all $r\in[1,\infty]$.
In this subsection, we establish the necessary conditions for Theorem 1.1. The condition $0\leq\frac{1}{q}$ is trivial, and the condition $\frac{1}{q}\leq\frac{1}{p}$ is a standard necessity following from a theorem of Hörmander [19].
We now demonstrate that, if
holds, then $(\frac{1}{p},\frac{1}{q})$ satisfies
$1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)\geq 0.$
The proof is similar to [31]. Specifically, it suffices to show
where the implicit constant is independent of $i$, and $\{t_{i}\}_{i=1}^{\infty}$ strictly decreasing and converging to 0. Indeed, let $S_{t_{i}}:=[-t_{i},t_{i}]\times[-|\gamma(t_{i})|,|\gamma(t_{i})|]$. As shown in [31] that
$\delta t_{i}\chi_{\frac{S_{t_{i}}}{2}}(x)\leq T\chi_{S_{t_{i}}}(x,u)$
To obtain the other necessary conditions for Theorem 1.1, we also require the following proposition, which is based on Lemmas 3.2, 3.3 and 3.4.